Suppose that one solution contains 40% alcohol and another solution contains 80% alcohol. How many liters of each solution should be mixed to make 8.5 liters of a 70% alcohol solution?
Found 2 solutions by ikleyn, amalm06:
Answer by ikleyn(52781)

.
Similar problem was solved in the lesson  Solving typical word problems on mixtures for solutions  in this site  (Problem 1).

Read this solution attentively and then substitute your data.

Good luck!


Answer by amalm06(224)

Use the method of alligation.

Let x denote the 40% solution and y denote the 80% solution.

Then 70-40=30 and 80-70=10

So that x/y=1/3

Find the quantity of x: x=(1/4)(8.5)=2.125 L (Answer)

y=8.5-2.125=6.375 L (Answer)